# Patch-Depth Fusion: Dichotomous Image Segmentation via Fine-Grained Patch Strategy and Depth Integrity-Prior

### v1

@@ Appendix E More Qualitative Evaluations
To better compare the differences between our method and other methods, we compare our VDPINet with other well-known task-related models. These models include CNN-based models such as IS-Net [1], FP-DIS [9], UDUN [10] and InSPyReNet [11], Transformer-based models like BiRefNet [12] and MVANet [2], and diffusion-based models GenPercept [13] and DiffDIS [3]. In addition, we also compare with the recent RGB-D SOD methods such as MAGNet [16] and CPNet [17], and RGB-D COD methods including DACOD [7] and RISNet [18] after they are trained on the DIS task. As shown in the Fig. A5, our method yields more accurate results for more complex scenarios, providing clearer foreground masks. Moreover, it still produces precise outputs for complex images. It can be seen that our method yields better results for transparent glass, regions with complex colors, and regions with similar colors.

### v2

@@ Appendix E More Qualitative Evaluations
To better compare the differences between our method and other methods, we compare our PDFNet with other well-known task-related models. These models include CNN-based models such as IS-Net [1], FP-DIS [9], UDUN [10] and InSPyReNet [11], Transformer-based models like BiRefNet [12] and MVANet [2], and diffusion-based models GenPercept [13] and DiffDIS [3]. In addition, we also compare with the recent RGB-D SOD methods such as MAGNet [16] and CPNet [17], and RGB-D COD methods including DACOD [7] and RISNet [18] after they are trained on the DIS task. As shown in the Fig. A5, our method yields more accurate results for more complex scenarios, providing clearer foreground masks. Moreover, it still produces precise outputs for complex images. It can be seen that our method yields better results for transparent glass, regions with complex colors, and regions with similar colors.